OpenStudy (anonymous):

Simultaneous equations: 3x+2y=7 2x-3y=-4 How do you solve this?

OpenStudy (paki):

here you need to find out the values of "x" and "y"...

OpenStudy (anonymous):

Yes, I know that

OpenStudy (anonymous):

@Hero Can you help?

OpenStudy (shinalcantara):

you can do elimination method. 3x+2y=7 --- equation 1 2x-3y=-4 ----equation 2 multiply equation 1 with 2 and equation 2 by 3 (3x+2y=7) * 2 (2x-3y=-4) * 3 (from eqn 1) 6x + 4y = 14 ----equation 1' (from eqn 2) 6x - 9y = -12 ----equation 2' subtract equations 1' & 2' 6x + 4y = 14 - 6x - 9y = -12 13y = 26 y = 2 substitute to equation 1 or 2 let's use eqn 1 3x+2y=7 3x+2(2)=7 3x + 4 = 7 3x = 7 - 4 3x = 3 x = 1

OpenStudy (anonymous):

@shinalcantara You're such a lifesaver. I did the first part, which was multiplying them to eliminate x, but I wasn't sure what to do next. Thank you so much!

OpenStudy (texaschic101):

you can even check your answers.. when x = 1 and y = 2 3x + 2y = 7 3(1) + 2(2) = 7 3 + 4 = 7 7 = 7 (correct)

OpenStudy (shinalcantara):

you can multiply any number but make sure you're gonna multiply it to the whole equation.. and you can do either subtraction or addition to the two equations which you do the multiplication as long as a variable can be eliminated.. try to practice other problems so you might understand.. :)

OpenStudy (shinalcantara):

yw :)

OpenStudy (anonymous):

@shinalcantara Thanks for the advice! You're really helpful!

OpenStudy (anonymous):

You said he wasn't a lifesaver